OpenStudy (gudden):

An alloy needs to contain between 46% copper and 50% copper. Find the least and greatest amounts of a 60% copper alloy that should be mixed with a 40% copper alloy in order to end up with thirty pounds of an alloy containing an allowable percentage of copper.

OpenStudy (taylor<3srin):

Let x = amount of 60% copper alloy then 30-x = amount of 40% copper alloy Least amount: .60x + .40(30-x) = .46(30) .60x + 12 - .40x = 13.8 .20x + 12 = 13.8 .20x = 1.8 x = 9 pounds . Greatest amount: .60x + .40(30-x) = .50(30) .60x + 12 - .40x = 15 .20x + 12 = 15 .20x = 3

OpenStudy (gudden):

thank you...

OpenStudy (taylor<3srin):

Close question ;)